# Bayesian Experimental Design for Model Discrepancy Calibration: An Auto-Differentiable Ensemble Kalman Inversion Approach

### v1

@@ 3 Numerical Results
To demonstrate the performance of our proposed method, we study the contaminant source inversion problem, which is a classical example for BED and has been previously studied in [96, 16]. The source inversion problem takes concentration measurements in a flow field governed by a convection-diffusion equation and then infers the plume source location. More specifically, the contaminant concentration $\mathbf{u}$ at a two-dimensional spatial location $\mathbf{z}=\{z_{x},z_{y}\}$ and time $t$ is governed by the following equation:
@@
where $\theta_{x}$ and $\theta_{y}$ denote the source location, and $\theta_{h}$ and $\theta_{s}$ represent the source width and source strength. The initial condition is $\mathbf{u}(\mathbf{z},0;\bm{\theta})=0$, and a homogeneous Neumann boundary condition is imposed for all sides of the square domain.
In this work, the parameters in the source term are set as $\{\theta_{x},\theta_{y},\theta_{h},\theta_{s}\}={0.25,0.25,2,0.05}$ and convection velocity set as $v_{x}=v_{y}=50t$, and Fig. 2 presents the system state $\mathbf{u}$ calculated at $t=0.03,0.06,0.09,0.12,0.15$ time units, which illustrates how the source affects the concentration value at different locations across the domain through the convection and diffusion with the time evolution.
Figure 2: Concentration value at different times in a convection-diffusion field. The numerical simulation is performed in a larger domain ($[-2,3]^{2}$) and presented in a smaller one ($[0,1]^{2}$) to emphasize the areas close to the source location.
The design $\mathbf{d}=\{d_{x},d_{y},d_{t}\}$ in this problem refers to the spatiotemporal coordinate to measure the concentration value. More specifically, each design involves the measurement of just one point in the domain. The spatial coordinates of the initial design start at $\{d_{x}^{0},d_{y}^{0}\}=(0.5,0.5)$ and will gradually move around to other locations. The difference of spatial coordinates between two consecutive designs $(|d_{x}^{i+1}-d_{x}^{i}|,|d_{y}^{i+1}-d_{y}^{i}|)$ is constrained by the interval $[-0.2,0.2]^{2}$ for physical parameters. The temporal coordinates of the measurement depend on different cases in general. At each time step, the posterior distribution from the previous stage serves as the prior for a new stage of BED as illustrated in Fig. 1. The current optimal design also serves as the starting point for the next stage of BED. In all numerical examples, the physics-based unknown parameters $\bm{\theta}_{\mathcal{G}}$ are the location $(\theta_{x},\theta_{y})$ of the source. The key motivations and findings of numerical results are summarized below:
@@
    We investigate the scalability of the proposed hybrid framework for BED with respect to the ensemble size and the AD-EKI iterations. The numerical results confirm the discussions about the peak memory and the time cost in Sections 2.3 and 2.4. More detailed results can be found in Section 3.3.
@@ 3.1 Parametric Model Error
The measurements are taken at $0.030,0.035,0.040,0.045,0.050,0.055$ time units as stages 1-6 of sequential BED. At the first stage, we employ a uniform prior distribution of $\{\theta_{x},\theta_{y}\}$ over the range $[0,1]^{2}$, discretized into a $51\times 51$ grid of points, and set an initial value of $\theta_{s}=3$. In each stage, standard BED is performed to infer the posterior distribution of $\{\theta_{x},\theta_{y}\}$, and the optimal design for $\{\theta_{x},\theta_{y}\}$ is determined accordingly based on the information gain introduced in Section 2.1. At each stage (except for stage 1) after the distribution of $\{\theta_{x},\theta_{y}\}$ is updated, optimal design for $\theta_{s}$ is determined based on the information gain estimated by AD-EKI, which was introduced in Section 2.3, to guide the data acquisition from the true system and facilitate the calibration of the model error $\theta_{s}$. The ensemble size is set as $30$, and the Gaussian distribution for the initial ensemble is assumed with the mean as the current value of $\theta_{s}$ and the variance as $1$.
@@
where $\lambda_{1}$ and $\lambda_{2}$ are the eigenvalues of the posterior covariance matrix. Compared to Fig. 3, Fig 4(a) presents the performance of the inferred posterior distribution in a more quantitative way. The blue solid line represents the evolution of the distance metric $D$ across different stages, while the shaded region indicates the corresponding magnitude of uncertainty $\sigma_{\text{eq}}$. The overall decrease in distance confirms the improvement in parameter estimation accuracy. However, a slight increase at stage 2 can be observed, which corresponds to the transition of the posterior distribution from a bimodal to an unimodal shape, leading to a temporary shift in the MAP estimate.
@@
Figure 3: Posterior results of physical parameter $\{\theta_{x},\theta_{y}\}$ via the hybrid approach on parametric error. The $\{\theta_{x},\theta_{y}\}$ space is $[0,1]^{2}$ as before, with a zoomed-in view $[0.2,0.8]\times[0.0,0.6]$ to highlight detailed behaviors.
We also examine the results of corrected parametric model error $\theta_{s}$. As shown in Fig. 4(b), the estimated value of $\theta_{s}$ gradually converges to the true value $\theta^{\dagger}_{s}=2$ over correction iterations. This result confirms that the proposed framework effectively adjusts for model discrepancy, ensuring that the parametric error value aligns closely with the true underlying model. Compared to a previous work [16], where the calibration of the model error parameter used optimal data from BED for physical parameters (i.e., source location in this example) and occasionally exhibited deviations from the true value, the estimation of model error parameter remains more robust here. This suggests that the optimal experimental design based on AD-EKI effectively identifies informative data directly for the model error parameter, ensuring a robust performance in the calibration of the error parameter $\theta_{s}$, which in turn benefits the inference of the physical parameters $\{\theta_{x},\theta_{y}\}$.
To further evaluate how the proposed framework enabled by AD-EKI distinguishes between designs of varying informativeness, we study the evolution of an AD-EKI approximated KL divergence during the first stage of sequential BED. A total of $70$ design updates were performed during this stage. Figure 5(a) presents three designs: (i) the initial design, (ii) an intermediate ($10$-th) design, and (iii) the final design, selected along the search trajectory for the optimal design. The AD-EKI approximated KL-divergence in Fig. 5(b) is based on the average results obtained from 50 independent runs with randomly initialized ensembles. For each design, the increasing KL divergence along the iteration steps reflects the nature of EKI as an iterative updating process, where information from the data is progressively incorporated into the updated ensemble. Moreover, the differences between the three curves at each iteration step confirm that the final design leads to the largest approximated KL-divergence and thus is considered as most informative.
@@ 3.2 Structural Model Error
which differs from the exponentially decaying source term as defined in Eq. (21) for the true system. Without addressing this model structural error, the inference of source location through standard BED yields biased results. In this work, a neural network $\mathbf{NN}(z_{x},z_{y},\theta_{x},\theta_{y};\bm{\theta_{\text{NN}}})$ is employed to characterize the model structural error. We utilize a fully connected neural network consisting of 37 parameters, a configuration that introduces a higher-dimensional parameter space than the example of parametric model error in Section 3.1. The neural network takes a 2-dimensional input vector, representing the relative distance from a specific grid point to the source location. The output is a scalar value for this point, indicating the discrepancy between the true and modeled source term values. The fully connected network is designed with two hidden layers, each containing four neurons, and employs a tanh activation function for both layers. The measurements are taken at $0.030,0.035,0.040,0.045,0.050,0.055$ time units as stages 1-6 of sequential BED. We employ a uniform prior distribution over the range $[0,1]^{2}$ for $\{\theta_{x},\theta_{y}\}$, discretized into a $51\times 51$ grid, with the true value of the source location set at $\{0.25,0.25\}$. The values of $\{\theta_{s}^{\dagger},\theta_{h}^{\dagger}\}$ are set as $\{2,0.05\}$ in both the true system and the modeled one. The advection velocity is known and set as $v_{x}=v_{y}=50t$. The ensemble size for network parameters is set to $40$. Each element of $\bm{\theta}_{\text{NN}}$ is assumed to have a zero-mean Gaussian error with a variance of $0.09$ for the prior distribution.
@@
Figure 8(b) presents the information gain at each iteration of the AD-EKI for three design stages: (i) the initial design, (ii) an intermediate design, and (iii) the final design as shown in Fig. 8(a). In total, $64$ design updates were performed during the EKI process. Although the initial design optimization was set for $70$ steps, the process was automatically terminated at the $64$-$th$ step due to reaching the predefined design search boundaries, triggering the stopping criterion. The $15$-$th$ design has a physical location approximately halfway between the initial and final designs and is selected to illustrate intermediate results of information gain over the optimization. The information gain of the updated designs consistently exceeds that of the preceding ones at each EKI iteration, which confirms the effectiveness of the ensemble-based information gain as a criterion for guiding the design optimization procedure.
@@
where $\mathbf{u}$ is the solution field from the modeled system and $\mathbf{u}^{\dagger}$ is the one from the true system. We first present the total error across the entire domain (i.e., $[0,1]^{2}$), which provides a general evaluation of the correction performance. Additionally, we show the local error in a smaller region (i.e., ±$\pm 0.04$ in both x and y directions) near the design for the physical parameters in the current stage, as an accurate estimation in this region is crucial for improving the Bayesian updating of the physical parameters. Since the model passed to the next stage is immediately used for design optimization, we also evaluate the effect of model correction on the subsequent stage. Specifically, we use the corrected model to optimize the next-stage design (for the Bayesian inference of physical parameters, i.e., the source location in this example) and compute the local error near the optimal design, comparing it to the results obtained using the uncorrected model.

### v2

@@ 3 Numerical Results
To demonstrate the performance of our proposed method, we study the contaminant source inversion problem, which is a classical example for BED and has been previously studied in [huan_gradient-based_2014, shen_bayesian_2023, yang_active_2025]. The source inversion problem takes concentration measurements in a flow field governed by a convection-diffusion equation and then infers the plume source location. More specifically, the contaminant concentration $\mathbf{u}$ at a two-dimensional spatial location $\mathbf{z}=\{z_{x},z_{y}\}$ and time $t$ is governed by the following equation:
@@
where $\theta_{x}$ and $\theta_{y}$ denote the source location, and $\theta_{h}$ and $\theta_{s}$ represent the source width and source strength. The initial condition is $\mathbf{u}(\mathbf{z},0;\bm{\theta})=0$, and a homogeneous Neumann boundary condition is imposed for all sides of the square domain $z_{L},z_{R}]^{2}=[-2,3]^{2}$.
In this work, the parameters in the source term are set as $\{\theta_{x},\theta_{y},\theta_{h},\theta_{s}\}={0.25,0.25,2,0.05}$ and convection velocity set as $v_{x}=v_{y}=50t$ and Fig. 2 presents the system state $\mathbf{u}$ calculated at $t=0.03,0.06,0.09,0.12,0.15$ time units, which illustrates how the source affects the concentration value at different locations across the domain through the convection and diffusion with the time evolution. The physical parameter support domain is set as $0,1]^{2}$, which is smaller than the PDE solving domain to control computational cost and avoid boundary effect. Similar settings can also be found in [shen_bayesian_2023].
Figure 2: Concentration value at different times in a convection-diffusion field. The numerical simulation is performed in a larger domain ($2,3]^{2}$) and presented in a smaller one ($0,1]^{2}$) to emphasize the areas close to the source location.
The design $\mathbf{d}=\{d_{x},d_{y},d_{t}\}$ in this problem refers to the spatiotemporal coordinate to measure the concentration value. More specifically, each design involves the measurement of just one point in the domain. The spatial coordinates of the initial design start at $\{d_{x}^{0},d_{y}^{0}\}=(0.5,0.5)$ and will gradually move around to other locations. The difference of spatial coordinates between two consecutive designs $(|d_{x}^{i+1}-d_{x}^{i}|,|d_{y}^{i+1}-d_{y}^{i}|)$ is constrained by the interval $0.2,0.2]^{2}$ for physical parameters. The temporal coordinates of the measurement depend on different cases in general. At each time step, the posterior distribution from the previous stage serves as the prior for a new stage of BED as illustrated in Fig. 1. The current optimal design also serves as the starting point for the next stage of BED. In all numerical examples, the physics-based unknown parameters $\bm{\theta}_{\mathcal{G}}$ are the location $(\theta_{x},\theta_{y})$ of the source. The key motivations and findings of numerical results are summarized below:
@@
    We investigate the scalability of the proposed hybrid framework for BED with respect to the ensemble size and the AD-EKI iterations. The numerical results confirm the discussions about the peak memory and the time cost in Sections 2.3 and 2.4. More detailed results can be found in Section $3.$
@@ 3.1 Parametric Model Error
The measurements are taken at $0.030,0.035,0.040,0.045,0.050,0.055$ time units as stages 1-6 of sequential BED. At the first stage, we employ a uniform prior distribution of $\{\theta_{x},\theta_{y}\}$ over the range $0,1]^{2}$, discretized into a $51\times 51$ grid of points, and set an initial value of $\theta_{s}=3$. In each stage, standard BED is performed to infer the posterior distribution of $\{\theta_{x},\theta_{y}\}$, and the optimal design for $\{\theta_{x},\theta_{y}\}$ is determined accordingly based on the information gain introduced in Section 2.1. At each stage (except for stage 1) after the distribution of $\{\theta_{x},\theta_{y}\}$ is updated, optimal design for $\theta_{s}$ is determined based on the information gain estimated by AD-EKI, which was introduced in Section 2.3, to guide the data acquisition from the true system and facilitate the calibration of the model error $\theta_{s}$. The ensemble size is set as $30$, and the Gaussian distribution for the initial ensemble is assumed with the mean as the current value of $\theta_{s}$ and the variance as $1$.
@@
where $\lambda_{1}$ and $\lambda_{2}$ are the eigenvalues of the posterior covariance matrix. Compared to Fig. 3, Fig 4(a) presents the performance of the inferred posterior distribution in a more quantitative way. The blue solid line represents the evolution of the distance metric $D$ across different stages, while the shaded region indicates the corresponding magnitude of uncertainty $\sigma_{\text{eq}}$. The overall decrease in distance confirms the improvement in parameter estimation accuracy.
@@
Figure 3: Posterior results of physical parameter $\{\theta_{x},\theta_{y}\}$ via the hybrid approach on parametric error. The $\{\theta_{x},\theta_{y}\}$ space is $0,1]^{2}$ as before, with a zoomed-in view $0.2,0.8]\times[0.0,0.6]$ to highlight detailed behaviors.
We also examine the results of corrected parametric model error $\theta_{s}$. As shown in Fig. 4(b), the estimated value of $\theta_{s}$ gradually converges to the true value $\theta^{\dagger}_{s}=2$ over correction iterations. This result confirms that the proposed framework effectively adjusts for model discrepancy, ensuring that the parametric error value aligns closely with the true underlying model. Compared to a previous work [yang_active_2025], where the calibration of the model error parameter used optimal data from BED for physical parameters (i.e., source location in this example) and occasionally exhibited deviations from the true value, the estimation of model error parameter remains more robust here. This suggests that the optimal experimental design based on AD-EKI effectively identifies informative data directly for the model error parameter, ensuring a robust performance in the calibration of the error parameter $\theta_{s}$, which in turn benefits the inference of the physical parameters $\{\theta_{x},\theta_{y}\}$.
To further evaluate how the proposed framework enabled by AD-EKI distinguishes between designs of varying informativeness, we study the evolution of an AD-EKI approximated KL divergence during the first stage of sequential BED. A total of $70$ design updates were performed during this stage. Figure 5(a) presents three designs: (i) the initial design, (ii) an intermediate ($10$-th) design, and (iii) the final design, selected along the search trajectory for the optimal design. The AD-EKI approximated KL divergence in Fig. 5(b) is based on the average results obtained from 50 independent runs with randomly initialized ensembles. For each design, the increasing KL divergence along the iteration steps reflects the nature of EKI as an iterative updating process, where information from the data is progressively incorporated into the updated ensemble. Moreover, the differences between the three curves at each iteration step confirm that the final design leads to the largest approximated KL divergence and thus is considered as most informative.
@@ 3.2 Structural Model Error
which differs from the exponentially decaying source term as defined in Eq. (21) for the true system. Without addressing this model structural error, the inference of source location through standard BED yields biased results. In this work, a neural network $\mathbf{NN}(z_{x},z_{y},\theta_{x},\theta_{y};\bm{\theta_{\text{NN}}})$ is employed to characterize the model structural error. We utilize a fully connected neural network consisting of 37 parameters, a configuration that introduces a higher-dimensional parameter space than the example of parametric model error in Section 3.1. The neural network takes a 2-dimensional input vector, representing the relative distance from a specific grid point to the source location. The output is a scalar value for this point, indicating the discrepancy between the true and modeled source term values. The fully connected network is designed with two hidden layers, each containing four neurons, and employs a tanh activation function for both layers. The measurements are taken at $0.030,0.035,0.040,0.045,0.050,0.055$ time units as stages 1-6 of sequential BED. We employ a uniform prior distribution over the range $0,1]^{2}$ for $\{\theta_{x},\theta_{y}\}$, discretized into a $51\times 51$ grid, with the true value of the source location set at $\{0.25,0.25\}$. The values of $\{\theta_{s}^{\dagger},\theta_{h}^{\dagger}\}$ are set as $\{2,0.05\}$ in both the true system and the modeled one. The advection velocity is known and set as $v_{x}=v_{y}=50t$. The ensemble size for network parameters is set to $40$. Each element of $\bm{\theta}_{\text{NN}}$ is assumed to have a zero-mean Gaussian error with a variance of $0.09$ for the prior distribution.
@@
Figure 8(b) presents the information gain at each iteration of the AD-EKI for three design stages: (i) the initial design, (ii) an intermediate design, and (iii) the final design as shown in Fig. 8(a). In total, $64$ design updates were performed during the EKI process. Although the initial design optimization was set for $70$ steps, the process was automatically terminated at the $64$-thth step due to reaching the predefined design search boundaries, triggering the stopping criterion. The $15$-thth design has a physical location approximately halfway between the initial and final designs and is selected to illustrate intermediate results of information gain over the optimization. The information gain of the updated designs consistently exceeds that of the preceding ones at each EKI iteration, which confirms the effectiveness of the ensemble-based information gain as a criterion for guiding the design optimization procedure.
@@
where $\mathbf{u}$ is the solution field from the modeled system and $\mathbf{u}^{\dagger}$ is the one from the true system. We first present the total error across the entire domain (i.e., $0,1]^{2}$), which provides a general evaluation of the correction performance. Additionally, we show the local error in a smaller region (i.e., $\pm 0.04$ in both x and y directions) near the design for the physical parameters in the current stage, as an accurate estimation in this region is crucial for improving the Bayesian updating of the physical parameters. Since the model passed to the next stage is immediately used for design optimization, we also evaluate the effect of model correction on the subsequent stage. Specifically, we use the corrected model to optimize the next-stage design (for the Bayesian inference of physical parameters, i.e., the source location in this example) and compute the local error near the optimal design, comparing it to the results obtained using the uncorrected model.
@@ 3.3 Empirical Results on Scalability of AD-EKI
4 Discussion
A key consideration is the robustness of the ensemble-based KL estimation, which relies on Gaussian assumptions and a finite ensemble size in practice. The ensemble-based KL divergence approximation in Eq. (7) is indeed formally derived under Gaussian assumptions. However, numerous studies have demonstrated that this hybrid Monte Carlo/Kalman-gain approach remains remarkably robust even when the underlying posterior distribution is mildly non-Gaussian [evensen2003ensemble, roth2017ensemble, evensen2022data]. We acknowledge that performance can be degraded for strongly skewed or multimodal posteriors, yet the recent successes in applying ensemble Kalman methods to neural network parameter estimation [chen2019approximate, lopez2022training] suggest that much less complicated posteriors are often expected in real applications. Some research indicates that the posterior generated by EKI updates can become biased or eventually collapse into pure noise in later iterations [pensoneault2024efficient], and our early stopping strategy effectively ensures the posterior remains informative and effective for the specific task of guiding experimental design.
Beyond the Gaussian assumption, challenges related to small ensemble sizes also warrant consideration. Insufficient ensemble sizes restrict the sample covariance to a low-dimensional subspace of the full parameter space, which restricts Kalman updates to that subspace and thus cannot correct errors orthogonal to it, resulting in biased estimates, underestimation of uncertainty, and potential divergence or slowed convergence of the ensemble Kalman methods. Given this potential issue of using ensemble Kalman methods, various techniques have been developed to mitigate challenges arising from small ensembles, including covariance inflation [anderson1999monte], localization [gaspari1999construction, houtekamer1998data], dropout [liu_dropout_2024], and other covariance estimation methods [nino2015ensemble, ledoit2004well, vishny_high-dimensional_2024]. Most of those techniques can be straightforwardly integrated into our framework to further enhance its robustness against non-Gaussianity or small ensemble effects. As the core contribution of this paper is to propose a general framework for applying EKI to BED, we leave the detailed integration and evaluation of these specific enhancement techniques as a direction for future work.
One potential concern in BED is that the EIG can be highly sensitive to the choice of prior. More specifically, if the prior is misspecified, an EIG-based design may preferentially collect data that simply reinforce those incorrect beliefs, yielding little practical insight [dasgupta1991robust, muller2012measuring, go2022robust]. In our AD-EKI framework, this sensitivity is substantially reduced in both physical-parameter and network-parameter settings. For physical parameters, we begin with a broad range of uniform priors that are weakly informative and periodically recondition on an improved forward model, so that any bias in early posteriors is progressively mitigated as more accurate model corrections become available. More results can be found in B. For high-dimensional network parameters, ensemble Kalman methods feature a theoretical invariance, i.e., if the affine span of the initial ensemble covers the entire parameter space, then the posterior mean can, in principle, locate to any true parameter point, regardless of the prior’s high-density region. This implies that AD-EKI is theoretically insensitive to the specific shape of the network prior and depends only on subspace coverage. In practice, finite ensemble sampling introduces Monte Carlo errors, while AD-EKI allows resampling of the ensemble at each gradient update to mitigate these errors.
In terms of the scalability with respect to network size, i.e., the network parameter dimension $d_{\theta}$, AD-EKI requires $\mathcal{O}(J\,d_{\theta})$ memory, with $\mathcal{O}(d_{\theta}^{2})$ for standard EKI and $\mathcal{O}(J\,d_{\theta})$ for dropout-EKI. In terms of convergence, the ensemble-based gradient estimate in AD-EKI retains an $O(J^{-1/2})$ error bound, which is the same as EKI since the KL divergence between Gaussians is Lipschitz continuous in both mean and covariance. Rigorous proof can be found in the related work AD-EnKF [chen_autodifferentiable_2022]. Though this is explicitly independent of $d_{\theta}$, typically standard EKI requires a similar amount of $J$ as $d_{\theta}$ to achieve this convergence rate. Dropout, or other localized covariance estimation techniques, may relax this requirement. Stability of AD-EKI stems from using derivative-free covariance updates in the inner loop of this bilevel optimization problem and from limiting the differentiation graph depth to a modest number of ensemble iterations. When the network dimension is huge such that drawing enough ensemble members is infeasible, additional techniques such as covariance inflation, dropout-based ensemble diversification, or other methods that break the subspace constraint may be applied to improve the robustness and stability of ensemble Kalman methods, thus facilitating the efficient use AD-EKI in real-world large-scale problems.
@@ Appendix A Details of derivation for the gradient of ensemble-based KL divergence
Appendix B Sensitivity test to prior of physical parameters
As mentioned in the Discussion, prior sensitivity is one concern in BED [dasgupta1991robust, muller2012measuring, go2022robust]. In this section, we provide results with different priors of physical parameters to validate our argument about the insensitivity of our framework to priors.
In the numerical tests in Sections 3.1 and 3.2, we have already demonstrated that our method works well with uniform priors.
We further consider the misspecified, which is biased, priors in the parametric case. In particular, we consider four types of misspecified 2D Gaussian priors: (1) flat and broad, (2) peaked and centered near the true value, (3) peaked and centered far from the true value, and (4) initially peaked and biased, but subsequently flattened into a near-uniform distribution, as illustrated in Fig. 11. Specifically, we obtain case (4) by raising each density value of the case (3) prior to the $1/5$ power and then renormalizing so that the transformed distribution sums to one. The error correction trajectories (Fig. 12) and posterior distributions (Fig. 13) in cases (1) and (2) demonstrate that our method remains robust under misspecified but moderately inaccurate priors. These results indicate a low sensitivity to priors. Although the convergence rate is slow under strongly biased priors (case 3), flattening such priors into a broader distribution (case 4) restores the effectiveness of the AD-EKI framework. Flattening the prior essentially places less weight on the prior and more on the data, effectively transforming a biased prior into a near-uniform one, under which our method has already been shown to be effective.
These findings reinforce the practical robustness of our method under a wide range of prior assumptions, justifying its applicability in scenarios with uncertain or biased prior knowledge.
Figure 11: Different 2D Gaussian priors for the physical parameters. From left to right: (1) flat and broad prior, (2) peaked prior centered near the true value, (3) peaked prior centered far from the true value, and (4) initially peaked and biased prior flattened into a broader distribution.
Figure 12: Error parameter correction trajectory with different priors.
Figure 13: Posterior results of the AD-EKI framework under different physical priors. From top to bottom: (1) flat and broad prior, (2) peaked prior centered near the true value, (3) peaked prior centered far from the true value, and (4) peaked and biased prior flattened into a broader distribution.